# An open-source heuristic to reboot 2D nesting research

### v1

@@ 1 Introduction
2D irregular cutting and packing (C&P) – or nesting – problems involve fitting a set of smaller items into larger containers, where the shape of the items and/or containers can be irregular (non-rectangular). Such optimization problems arise in a wide range of industries such as garment manufacturing, furniture production, shipbuilding, printing, laser-cutting, metalworking, woodworking, foam-cutting, leather crafting, and microchip design. Designing effective algorithms to solve these NP-hard problems (Fowler et al., 1981) has been an active area of research since the 1960s (Art Jr, 1966).
Given the real-world relevance of 2D irregular C&P problems and the ongoing shift in ethos toward open-source practices (UNESCO, 2021), one would expect to find a vibrant ecosystem of publicly available implementations that respond to advances in the academic state of the art (SotA). However, despite more than a decade of stagnation – where there has been little to no improvement in achievable solution quality – the open-source landscape remains disconnected from academic progress and trails significantly behind in performance. We hypothesize that there are two interconnected factors primarily responsible for this situation.
@@
The second factor, we believe, which is leading to this stagnation is a lack of reproducibility. At the time of writing, not a single academically competitive algorithm for 2D irregular C&P problems has publicly available source code. Moreover, many methods fail to provide enough explanation to facilitate complete reproducibility.
This paper has two interrelated goals: (i) bridge the gap between academic and open-source algorithms and (ii) lower the barrier to entry for future researchers. When addressing these goals, we will rely on the recently introduced Collision Detection Engine (CDE) for 2D irregular C&P problems (Gardeyn et al., 2025), which can handle the geometric challenge. Our entire focus can therefore target the optimization challenge. On top of the CDE, we will design an open-source heuristic for the 2D irregular strip packing problem, arguably the most canonical 2D nesting problem. While we focus on strip packing, our broader ambition is to develop an algorithm that can generalize to other variants of 2D irregular C&P problems without much effort.
@@ 3 State of 2D irregular strip packing
Exact methods for the 2DISPP typically rely on mathematical programming formulations. Leao et al. (2020) provide a comprehensive survey of the broad range of exact algorithms for irregular C&P problems. Such approaches can generally only handle a small set of items, support few angles of rotation and struggle with complex shapes. Although more recent work – such as the mixed-integer formulation by Lastra-D´ıaz and Ortuño (2024) – demonstrates continued advances, exact approaches remain impractical for many academic instances, let alone complex real-world instances.
@@
Construction heuristics build solutions sequentially by placing one item at a time according to a predetermined ordering in addition to a placement rule. To improve solution quality, many methods embed a higher-level heuristic that explores different orderings and placement strategies. Representative work has been conducted by Oliveira et al. (2000); Gomes and Oliveira (2002); Burke et al. (2006); Bennell and Song (2010); Amaro Junior et al. (2017). These algorithms are generally straightforward to implement and reproducible provided that sufficient detail is included in the corresponding publication.
@@
Most methods relying on iterative improvement begin with a complete solution and repeatedly apply local modifications – such as shifting or swapping items – in an attempt to find new solutions of higher quality. Early contributions include Egeblad et al. (2007); Umetani et al. (2009); Leung et al. (2012). More recent advances by Elkeran (2013); Wang et al. (2017); Sato et al. (2019) represent the SotA for the 2DISPP. The common theme across these methods is that they tolerate temporary collisions between items. The strip gets converted into a container with fixed dimensions and, every time a feasible solution is found, its length is reduced further. This shrinking leads to new collisions, which the aforementioned methods attempt to gradually resolve in order to reach a feasible solution for the shorter strip.
While many of these iterative-improvements heuristics can achieve high-quality solutions, they tend to be more complicated than construction-based heuristics. Even when their algorithmic flow is presented clearly – as in Sato et al. (2019) – the absence of source code or sufficient implementation details prevents true reproducibility.
@@
Solid, robust and publicly available implementations to solve the 2DISPP are scarce. A notable exception is a tree-search heuristic by Fontan (2023). This heuristic has a unique approach to handling the geometric challenge which involves decomposing shapes into trapezoids. Despite representing a valuable open-source contribution, Fontan (2023)’s method is unrelated to the iterative-improvement methods and cannot consistently compete with the academic SotA.
@@ 4 A sequence of feasibility problems
Umetani et al. (2009) referred to this subproblem as the ‘Overlap Minimization Problem’, but we find this term somewhat misleading. Feasibility, not minimal overlap, is the true goal. Any solution with overlap is ultimately invalid, regardless of how minimal it may be. Therefore, a more appropriate name for this subproblem would probably be the ‘Overlap Elimination Problem’.
The idea of exploring infeasible regions has existed for a long time (Glover, 1989) and is a strategy commonly used in the irregular C&P literature. However, the reasons for its effectiveness in nesting problems are rarely elaborated upon. We will therefore take this opportunity to offer some intuition for why this conversion works so well.
@@
  • Search algorithm: A strategy which explores the solution space in order to incrementally resolve collisions.
@@ 5 Detecting collisions
The most elementary challenge involved in solving any feasibility problem is assessing whether or not a constraint is violated. Given the geometric challenge that accompanies nesting problems, checking the feasibility of a solution or even of a single placement in a way that is efficient, precise and robust is not trivial. The fundamental ways of dealing with this geometric challenge have been described by Bennell and Oliveira (2008). Currently, all state-of-the-art nesting algorithms rely on either no-fit polygons (NFPs) or a raster to handle the geometric challenge. However, these two approaches have serious limitations regarding robustness and precision, respectively.
A third approach mentioned by Bennell and Oliveira (2008) is trigonometry: a simple, precise, robust and general way of detecting collisions between two polygons. This approach was historically very sensitive to the total number of edges in the shapes involved, and therefore seldom used in practice.
Recently, Gardeyn et al. (2025) introduced a Collision Detection Engine (CDE) to provide a fast and reliable way for 2D nesting algorithms to assess feasibility. This CDE builds on trigonometric principles and includes a comprehensive suite of algorithmic enhancements to overcome the traditional limitations of such trigonometric approaches. The CDE is implemented in an open-source library called jagua-rs [^1] https://github.com/JeroenGar/jagua-rs. jagua-rs supports irregular shapes for both items and containers, continuous rotation and translation, and easily extends to problem variants containing additional spatial constraints.
@@ 6 Quantifying collisions
[…]
@@
Many existing approaches for the 2DISPP quantify the severity of a collision in some way. Imamichi et al. (2009) proposed using penetration depth, which denotes the smallest translation needed to separate two colliding entities. Instead of the exact penetration depth, they measure the minimum translation required along a single axis, such as horizontally or vertically. Leung et al. (2012), Elkeran (2013), Wang et al. (2017) and Sato et al. (2019) each employ some variant of penetration depth to quantify the severity of a collision. Regardless of the concrete implementation, these penetration-depth derivatives are essentially all one-dimensional metrics.
@@ 6.1 Area of overlap
A shape’s pole of inaccessibility is the point within its interior that lies furthest from the boundary. This point, along with its distance to the boundary, defines the largest inscribable circle in the shape. Figure 3 shows a sequence of such inscribed circles – referred to simply as poles from now on – for each shape. Gardeyn et al. (2025) describe the precise procedure to compute these sets of poles, extending the approach by Agafonkin (2016). The CDE leverages them to accelerate the collision detection process and lower its sensitivity to the complexity of the shapes. The set of poles of shape $S$ is represented by $P(S)$. Poles are generated once (in preprocessing) for every base shape and can be efficiently transformed in unison with any rigid transformation applied to the shape.
@@
$(p_{a},p_{b})\in P(S_{a})\times P(S_{b})\triangleright$ all pairs of poles
@@ 6.2 Decaying penetration depth
To address the aforementioned issue, we introduce a decaying version of the penetration depth between pairs of poles. When $\delta$ is smaller than a certain threshold $\varepsilon$, we switch to a hyperbolic decay function which asymptotically approaches 0 as the distance between the pair of poles increases. The decaying penetration depth $\delta^{\prime}$ is defined as follows:
@@
[…]
@@ 6.3 Shape-based penalty
The proxy provided by Algorithm 3 is a decent starting point for quantifying the severity of a collision. However, the area of overlap is not necessarily the only factor which influences how difficult a collision is to resolve. When quantifying collisions purely on the basis of the amount of overlap, we noticed that our search algorithm had a strong tendency to converge to situations with minor amounts of overlap between pairs of items with large and/or concave shapes. It seems that resolving such collisions is disproportionately more disruptive to the structure of the overall solution and therefore much harder compared to resolving collisions between smaller and/or more convex items.
@@
[…]
@@ 7 Navigating a continuous search space
All that now remains is to design a search algorithm which is able to gradually resolve collisions. One of the most fundamental components of any local search algorithm is the move operation, enabling the algorithm to transition from one solution to a neighboring one. Algorithm 5 introduces the local search move our approach will employ. Its core idea – repositioning every colliding item, one by one, in a randomized order – is shared by Umetani et al. (2009); Sato et al. (2019); Elkeran (2013).
@@
[…]
In 2D irregular C&P problems, items can be positioned in a continuous 2D space and sometimes also rotated in a continuous manner. This means it is far too time-consuming to evaluate a full neighborhood of all possible item positions. Previous approaches have handled this continuous search space in a variety of ways. Sato et al. (2019) discretize the search space into a (multi-resolution) grid. Meanwhile, approaches that rely on NFPs (Umetani et al., 2009; Elkeran, 2013) narrow the search space down to a set of positions along the edges of the union of all NFPs. From this (still continuous) set of positions, only a few specific points such as the vertices of the combined NFP are typically considered. All of these approaches effectively restrict the search space to a finite set of positions.
@@
[…]
@@
We therefore select some of the most promising samples $\in T_{div}\cup T_{foc}$ to be refined into their respective local optima. A sample is considered promising if it is among the best evaluated and its transformation is sufficiently different from the other samples already selected. Figure 7(d) visualizes this refinement process for three samples that are highlighted in lime, cyan and purple. The refinement procedure is inspired by the adaptive coordinate descent approach described by Loshchilov et al. (2011) and greedily shifts sampling in the direction of improving evaluation. Figure 7(e) shows the position eventually chosen to move the item to.
@@ 8.1 Guided local search
How to search for a new position was explained in Algorithm 6, but the function to evaluate samples was left abstract. A straightforward way to evaluate a sample would be to aggregate the severity of all collisions that would occur if the item were placed at the sampled position, akin to Algorithm 1. However, such a static evaluation function would cause the search to converge to a local optimum within a couple of iterations. Approaches proposed by Umetani et al. (2009) and Sato et al. (2019) avoid this issue by relying on Guided Local Search (GLS) (Voudouris et al., 2010): a metaheuristic which dynamically modifies the evaluation function to escape local optima and move towards promising regions of the search space.
@@
[…]
@@
[…]
@@ 8.2 Separation Procedure
Algorithm 9 separate($m_{\text{max}},n_{\text{max}}$)
@@
3:move_items_multi $\triangleright$ Alg. 10
4:update_weights $\triangleright$ Alg. 8
[…]
Algorithm 9 describes the complete procedure to solve the feasibility problem introduced in Section 4. The separation procedure starts from an infeasible situation with colliding items and attempts to resolve all of its collisions. At its core, the algorithm continually repositions items, updates the item-pair weights and replaces the incumbent solution $s^{*}$ whenever improvements are found with respect to Equation 8. Whenever feasibility is reached, the procedure stops and immediately returns the solution. Once a certain number of iterations without improvements $n_{\text{max}}$ is reached, we conclude one attempt. If the incumbent solution was never updated during this attempt, a strike is added. Conversely, if the attempt improved $s^{*}$, the strike counter is reset to 0. As long as the number of strikes is below a certain threshold $m_{\text{max}}$, the procedure restores the incumbent solution $s^{*}$ and starts a new attempt, continuing its quest for a feasible solution. Once the number of strikes exceeds $m_{\text{max}}$, the procedure halts the search and returns $s^{*}$, which will unfortunately still be infeasible.
The behavior we intend to generate through this procedure can be described as follows. GLS ensures that item pairs colliding across consecutive iterations quickly reach prohibitively large weights, especially pairs which are involved in the most severe collisions. By contrast, when a pair of items manages to resolve its collision, the weight slowly decays back down to 1. When a collision occurs in a region with low GLS weights, the items involved will often ‘transfer’ the collision to neighboring pairs. Observing this process over multiple iterations reveals that items effectively push and pull each other, constantly ‘fighting’ for their own space in the container. When the search inevitably stalls, some items start to accumulate significant weights with all their neighbors. This will cause them to jump away to an entirely different position in the container, effectively escaping the local optimum. Readers who are curious to see this behavior in action are referred to the repository (Section 10), which contains a simple visualizer.

### v2

@@ 1 Introduction
2D irregular cutting and packing (C&P) – or nesting – problems involve fitting a set of smaller items into larger containers, where the shape of the items and/or containers can be irregular (non-rectangular). Such optimization problems arise in a wide range of industries such as garment manufacturing, furniture production, shipbuilding, printing, laser-cutting, metalworking, woodworking, foam-cutting, leather crafting, and microchip design. Designing effective algorithms to solve these NP-hard problems (fowler1981optimal) has been an active area of research since the 1960s (art1966approach).
Given the real-world relevance of 2D irregular C&P problems and the ongoing shift in ethos toward open-source practices (unesco2021openscience), one would expect to find a vibrant ecosystem of publicly available implementations that respond to advances in the academic state of the art (SotA). However, despite more than a decade of stagnation – where there has been little to no improvement in achievable solution quality – the open-source landscape remains disconnected from academic progress and trails significantly behind in performance. We hypothesize that there are two interconnected factors primarily responsible for this situation.
@@
The second factor, we believe, that is leading to this stagnation is a lack of reproducibility. At the time of writing, not a single academically competitive algorithm for 2D irregular C&P problems has publicly available source code. Moreover, many methods fail to provide enough explanation to facilitate complete reproducibility.
This paper has two interrelated goals: (i) bridge the gap between academic and open-source algorithms and (ii) lower the barrier to entry for future researchers. When addressing these goals, we will rely on the recently introduced Collision Detection Engine (CDE) for 2D irregular C&P problems (gardeyn2025decoupling), which can handle the geometric challenge. Our entire focus can therefore target the optimization challenge. On top of the CDE, we will design an open-source heuristic for the 2D irregular strip packing problem, arguably the most canonical 2D nesting problem. While we focus on strip packing, our broader ambition is to develop an algorithm that can generalize to other variants of 2D irregular C&P problems without much effort.
@@ 3 State of 2D irregular strip packing
Exact methods for the 2DISPP typically rely on mathematical programming formulations. leao2020irregular provide a comprehensive survey of the broad range of exact algorithms for irregular C&P problems. Such approaches can generally only handle a small set of items, support few angles of rotation and struggle with complex shapes. Although more recent work – such as the mixed-integer formulation by lastra2024mixed – demonstrates continued advances, exact approaches remain impractical for many academic instances, let alone complex real-world instances.
@@
Construction heuristics build solutions sequentially by placing one item at a time according to a predetermined ordering in addition to a placement rule. To improve solution quality, many methods embed a higher-level heuristic that explores different orderings and placement strategies. Representative work has been conducted by oliveira2000topos; gomes2002two; burke2006new; bennell2010beam; amaro2017parallel. These algorithms are generally straightforward to implement and reproducible provided that sufficient detail is included in the corresponding publication.
@@
Most methods relying on iterative improvement begin with a complete solution and repeatedly apply local modifications – such as shifting or swapping items – in an attempt to find new solutions of higher quality. Early contributions include egeblad2007fast; umetani2009solving; leung2012extended. More recent advances by elkeran2013new; wang2017flexible; sato2019raster represent the SotA for the 2DISPP. The common theme across these methods is that they tolerate temporary collisions between items. The strip gets converted into a container with fixed dimensions and, every time a feasible solution is found, its length is reduced further. This shrinking leads to new collisions, which the aforementioned methods attempt to gradually resolve in order to reach a feasible solution for the shorter strip.
While many of these iterative-improvements heuristics can achieve high-quality solutions, they tend to be more complicated than construction-based heuristics. Even when their algorithmic flow is presented clearly – as in sato2019raster – the absence of source code or sufficient implementation details prevents true reproducibility.
@@
Solid, robust and publicly available implementations to solve the 2DISPP are scarce. A notable exception is a tree-search heuristic by fontan. This heuristic has a unique approach to handling the geometric challenge that involves decomposing shapes into trapezoids. Despite representing a valuable open-source contribution, fontan’s method is unrelated to the iterative-improvement methods and cannot consistently compete with the academic SotA.
@@ 4 A sequence of feasibility problems
umetani2009solving referred to this subproblem as the ‘Overlap Minimization Problem’, but we find this term somewhat misleading. Feasibility, not minimal overlap, is the true goal. Any solution with overlap is ultimately invalid, regardless of how small it may be. Additionally, reaching a feasible solution does not entail strictly minimizing overlap, there may be additional factors that aid the search in its goal to eliminate overlap. Therefore, a more appropriate name for this subproblem would probably be the ‘Overlap Elimination Problem’.
The idea of exploring infeasible regions has existed for a long time (glover1989tabu) and is a strategy commonly used in the irregular C&P literature. However, the reasons for its effectiveness in nesting problems are rarely elaborated upon. We will therefore take this opportunity to offer some intuition for why this conversion works so well.
@@
  • 2.
    Search algorithm: A strategy that explores the solution space in order to incrementally resolve collisions.
  • 3.
@@ 5 Detecting collisions
The most elementary challenge involved in solving any feasibility problem is assessing whether or not a constraint is violated. Given the geometric challenge that accompanies nesting problems, checking the feasibility of a solution or even of a single placement in a way that is efficient, precise and robust is not trivial. The fundamental ways of dealing with this geometric challenge have been described by bennell2008geometry. Currently, all state-of-the-art nesting algorithms rely on either no-fit polygons (NFPs) or a raster to handle the geometric challenge. However, these two approaches have serious limitations regarding robustness and precision, respectively.
A third approach mentioned by bennell2008geometry is trigonometry: a simple, precise, robust and general way of detecting collisions between two polygons. This approach was historically very sensitive to the total number of edges in the shapes involved, and therefore seldom used in practice.
Recently, gardeyn2025decoupling introduced a Collision Detection Engine (CDE) to provide a fast and reliable way for 2D nesting algorithms to assess feasibility. This CDE builds on trigonometric principles and includes a comprehensive suite of algorithmic enhancements to overcome the traditional limitations of such trigonometric approaches. The CDE is implemented in an open-source library called jagua-rs [^1] https://github.com/JeroenGar/jagua-rs. jagua-rs supports irregular shapes for both items and containers, continuous rotation and translation, and easily extends to problem variants containing additional spatial constraints.
@@ 6 Quantifying collisions
1: if $a\in\texttt{jagua-rs::collisions}(S_{b})$ then
2:   return $\texttt{quantify\_collision}(S_{a},S_{b})$ $\triangleright$ Alg. 4
3: else
4:   return 0
@@
Many existing approaches for the 2DISPP quantify the severity of a collision in some way. imamichi2009iterated proposed using penetration depth, which denotes the smallest translation needed to separate two colliding entities. Instead of the exact penetration depth, they measure the minimum translation required along a single axis, such as horizontally or vertically. leung2012extended, elkeran2013new, wang2017flexible and sato2019raster each employ some variant of penetration depth to quantify the severity of a collision. Regardless of the concrete implementation, these penetration-depth derivatives are essentially all one-dimensional metrics.
@@ 6.1 Area of overlap
A shape’s pole of inaccessibility is the point within its interior that lies furthest from the boundary. This point, along with its distance to the boundary, defines the largest inscribable circle in the shape. Figure 3 shows a sequence of such inscribed circles – referred to simply as poles from now on – for each shape. gardeyn2025decoupling describe the precise procedure to compute these sets of poles, extending the approach by Agafonkin_Polylabel_a_fast_2016. The CDE leverages them to accelerate the collision detection process and lower its sensitivity to the complexity of the shapes. The set of poles of shape $S$ is represented by $P(S)$. Poles are generated once (in preprocessing) for every base shape and can be efficiently transformed in unison with any rigid transformation applied to the shape.
@@
1: for $(p_{a},p_{b})\in P(S_{a})\times P(S_{b})$ do $\triangleright$ all pairs of poles
2:   $\delta\leftarrow\textsl{penetration depth}(p_{a},p_{b})$
3:   if $\delta>0$ then
@@ 6.2 Decaying penetration depth
To address the aforementioned issue, we introduce a decaying version of the penetration depth between pairs of poles. When $\delta$ is smaller than a certain threshold $\varepsilon$, we switch to a hyperbolic decay function that asymptotically approaches 0 as the distance between the pair of poles increases. The decaying penetration depth $\delta^{\prime}$ is defined as follows:
@@
2: for $(p_{a},p_{b})\in P(S_{a})\times P(S_{b})$ do
3:   $\delta\leftarrow\textsl{penetration depth}(p_{a},p_{b})$
4:   if $\delta>\varepsilon$ then
5:    $\delta^{\prime}\leftarrow\delta$
6:   else
7:    $\delta^{\prime}\leftarrow\varepsilon^{2}/(-\delta+2\varepsilon)$
8:   $\alpha\leftarrow\alpha+\delta^{\prime}\cdot\text{min}\{\diameter(p_{a}),\diameter(p_{b})\}$
9: return $\alpha$
@@ 6.3 Shape-based penalty
The proxy provided by Algorithm 3 is a decent starting point for quantifying the severity of a collision. However, the area of overlap is not necessarily the only factor that influences how difficult a collision is to resolve. When quantifying collisions purely on the basis of the amount of overlap, we noticed that our search algorithm had a strong tendency to converge to situations with minor amounts of overlap between pairs of items with large and/or concave shapes. It seems that resolving such collisions is disproportionately more disruptive to the structure of the overall solution and therefore much harder compared to resolving collisions between smaller and/or more convex items.
@@
2: $\lambda_{ab}\leftarrow\sqrt{\lambda_{a}\cdot\lambda_{b}}$ $\triangleright$ Eq. 7
3: return $\sqrt{\alpha}\cdot\lambda_{ab}$
@@ 7 Navigating a continuous search space
All that now remains is to design a search algorithm that is able to gradually resolve collisions. One of the most fundamental components of any local search algorithm is the move operation, enabling the algorithm to transition from one solution to a neighboring one. Algorithm 5 introduces the local search move our approach will employ. Its core idea – repositioning every colliding item, one by one, in a randomized order – is shared by umetani2009solving; sato2019raster; elkeran2013new.
@@
1: $I_{c}\leftarrow\{i\in I:|\texttt{jagua-rs::collisions}(S_{i})|>1\}$ $\triangleright$ colliding items
2: for $i\in{I_{c}}$ in a random order do
3:   $t\leftarrow$ search_position(ii) $\triangleright$ Alg. 6
4:   jagua-rs::move_item($i,t$)
In 2D irregular C&P problems, items can be positioned in a continuous 2D space and sometimes also rotated in a continuous manner. This means it is far too time-consuming to evaluate a full neighborhood of all possible item positions. Previous approaches have handled this continuous search space in a variety of ways. sato2019raster discretize the search space into a (multi-resolution) grid. Meanwhile, approaches that rely on NFPs (umetani2009solving; elkeran2013new) narrow the search space down to a set of positions along the edges of the union of all NFPs. From this (still continuous) set of positions, only a few specific points such as the vertices of the combined NFP are typically considered. All of these approaches effectively restrict the search space to a finite set of positions.
@@
1: $f(i,t)\mapsto e\ \leftarrow$ evaluate_sample(i,t)(i,t) $\triangleright$ Alg. 7
2: $t^{*}\leftarrow\texttt{perform\_sampling}(i,f(i,t)\mapsto e)$ $\triangleright$ Fig. 7
3: return $t^{*}$
@@
We therefore select some of the most promising samples $\in T_{div}\cup T_{foc}$ to be refined into their respective local optima. A sample is considered promising if it is among the best evaluated and its transformation is sufficiently different from the other samples already selected. Figure 7(d) visualizes this refinement process for three samples that are highlighted in lime, cyan and purple. The refinement procedure is inspired by the adaptive coordinate descent approach described by loshchilov2011adaptive and greedily shifts sampling in the direction of improving evaluation. Figure 7(e) shows the position eventually chosen to move the item to.
@@ 8.1 Guided local search
How to search for a new position was explained in Algorithm 6, but the function to evaluate samples was left abstract. A straightforward way to evaluate a sample would be to aggregate the severity of all collisions that would occur if the item were placed at the sampled position, akin to Algorithm 1. However, such a static evaluation function would cause the search to converge to a local optimum within a couple of iterations. Approaches proposed by umetani2009solving and sato2019raster avoid this issue by relying on Guided Local Search (GLS) (voudouris2010guided): a metaheuristic that dynamically modifies the evaluation function to escape local optima and move towards promising regions of the search space.
@@
1: $C\leftarrow\texttt{jagua-rs::collisions}(t(S_{i}))$
2: for $c\in C\setminus\{i\}$ do $\triangleright$ ignore item $i$ itself
3:   $w_{ic}\leftarrow\text{weight of item-pair }(i,c)$
4:   $e\leftarrow e+w_{ic}\cdot\texttt{quantify\_collision}(t(S_{i}),S_{c})$ $\triangleright$ Alg. 4
5: return $e$
@@
2: for $a,b\in I:a\neq b$ do $\triangleright$ all pairs of items
3:   $e\leftarrow$ evaluate_item_pair(a,b)(a,b) $\triangleright$ Alg. 1
4:   if $e>0$ then
5:    $m\leftarrow\texttt{M}_{l}+(\texttt{M}_{u}-\texttt{M}_{l})\cdot(e/e_{\text{max}})$
6:   else
7:    $m\leftarrow\texttt{M}_{d}$
@@ 8.2 Separation Procedure
Algorithm 9 separate($k_{\text{max}},n_{\text{max}}$)
@@
2: while $k<k_{\text{max}}$ and $e^{*}>0$ do
3:   restore to $s^{*}$} ; $s_{\text{init}}\leftarrow s^{*}$ ; $n\leftarrow 0$
4:   while $n<n_{\text{max}}$ and $e^{*}>0$ do $\triangleright$ an attempt
5:    move_items_multi $\triangleright$ Alg. 10
6:    update_weights $\triangleright$ Alg. 8
@@
8:    if $e<e^{*}$ then
9:      $s^{*}\leftarrow$ current solution ; $e^{*}\leftarrow e$
10:      $n\leftarrow 0$
11:   $k\leftarrow k+1$ $\triangleright$ add a strike
12:   if $s^{*}\neq s_{\text{init}}$ then
13:    $k\leftarrow 0$ $\triangleright$ reset strikes
14: return $s^{*}$
Algorithm 9 describes the complete procedure to solve the feasibility problem introduced in Section 4. The separation procedure starts from an infeasible situation with colliding items and attempts to resolve all of its collisions. At its core, the algorithm continually repositions items, updates the item-pair weights and replaces the incumbent solution $s^{*}$ whenever improvements are found with respect to Equation 8. Whenever feasibility is reached, the procedure stops and immediately returns the solution. Once a certain number of iterations without improvements $n_{\text{max}}$ is reached, we conclude one attempt. If the incumbent solution was never updated during this attempt, a strike is added. Conversely, if the attempt improved $s^{*}$, the strike counter $k$ is reset to 0. As long as the number of strikes is below a certain threshold $k_{\text{max}}$, the procedure restores the incumbent solution $s^{*}$ and starts a new attempt, continuing its quest for a feasible solution. Once the number of strikes exceeds $k_{\text{max}}$, the procedure halts the search and returns $s^{*}$, which will unfortunately still be infeasible.
The behavior we intend to generate through this procedure can be described as follows. GLS ensures that item pairs colliding across consecutive iterations quickly reach prohibitively large weights, especially pairs that are involved in the most severe collisions. By contrast, when a pair of items manages to resolve its collision, the weight slowly decays back down to 1. When a collision occurs in a region with low GLS weights, the items involved will often ‘transfer’ the collision to neighboring pairs. Observing this process over multiple iterations reveals that items effectively push and pull each other, constantly ‘fighting’ for their own space in the container. When the search inevitably stalls, some items start to accumulate significant weights with all their neighbors. This will cause them to jump away to an entirely different position in the container, effectively escaping the local optimum. Readers who are curious to see this behavior in action are referred to the repository (Section 10), which contains a simple visualizer.